# The frustrated Ising model on the honeycomb lattice: Metastability and universality

### v1

@@ I Introduction
More recently, the same model on the honeycomb lattice has attracted some attention [18, 19, 20, 21, 22, 23, 24]. Unlike in non-frustrated spin systems, the lattice geometry plays a vital role for the level of frustration present, and therefore affects the macroscopic behavior of the system. This is illustrated by the textbook example of the antiferromagnetic nn Ising model on square and triangular lattices, respectively. The former maps to the ferromagnetic Ising model by a simple transformation, whereas the latter remains disordered at all temperatures and has a nonzero ground-state entropy per spin [25, *Wannier1973Errata]. In fact, precisely this example illustrates the reason why the behavior of the $J_{1}$-$J_{2}$ Ising model is expected to be different on the honeycomb lattice: On the square lattice, the nnn bonds form square lattices, whereas on a honeycomb lattice they form two triangular lattices. Thus, different phenomena may be anticipated for this case, particularly for strong antiferromagnetic nnn interactions.
@@
The structure of the ground states in the honeycomb system depends on the relative strength of $J_{1}$ and $J_{2}$ [27]. For $J_{2}>-|J_{1}|/4$, depending on the sign of $J_{1}$, the system either has a ferromagnetically ordered or a Néel-ordered ground state, whereas for $J_{2}<-|J_{1}|/4$ the ground-state manifold is highly degenerate. As is the case on the square lattice, a simple transformation maps negative $J_{1}$ to positive ones (and vice-versa). Only quite recently the model was first studied for non-zero temperatures by Bobák et al. [18]. For $J_{2}>-|J_{1}|/4$, using an effective-field theory (EFT) of clusters of different sizes they find evidence of a tricritical point separating a region of second-order transitions for $0.1\lesssim J_{2}/|J_{1}|\leq 0$ from one with first-order transitions for $0.25<J_{2}/|J_{1}|\lesssim-0.1$. One of the authors of Ref. [18] subsequently carried out a Monte Carlo (MC) simulation study [20] in which no tricritical point was observed. Instead it showed that the transition remains of second order and within the Ising universality class at least for $J_{2}/|J_{1}|\geq-0.2$. For even lower values of $J_{2}$ simulations showed hysteresis upon cooling and heating owed to a dramatic increase of the autocorrelation times. This increased range of the second-order regime as compared to that of Ref. [18] is supported by a cluster mean-field study [21] which finds only second-order signals for the entire range of $0.25<J_{2}/|J_{1}|\leq 0$. We recently presented some Monte Carlo results of the scaling of the partition function zeros [24], which indicated that the system remains in the Ising universality class at least for $J_{2}/|J_{1}|\geq-0.22$. For $J_{2}/|J_{1}|<-1/4$, on the other hand, it appears unclear whether there is a phase transition at all. Specific-heat peaks are suggestive of some sort of transition [28], and due to the extremely slow relaxation it has been argued that the system shows signs of spin-glass-like ordering [19]. Results of machine learning studies [22, 23] were interpreted as suggestive of two distinct phases seen for small system sizes, hence indicative of the existence of a phase transition.
@@
In the present work, however, we focus on the range in $J_{2}$ with long-range order at $T=0$, i.e., $J_{2}/|J_{1}|\in(-1/4,0]$. As one approaches $J_{2}\searrow-0.25|J_{1}|$, standard Monte Carlo techniques such as the Metropolis algorithm are found to be increasingly inefficient due to two effects: First, the introduction of the antiferromagnetic nnn interaction gives rise to local energy minima that simulations tend to get trapped in. Second, as $J_{2}$ is lowered the transition temperature decreases, which translates into low acceptance rates that are further reducing the efficacy of the Metropolis algorithm. For the present study we use the population annealing [29] framework which is well-suited for systems with rough free-energy landscapes [30], and which has excellent parallel efficiency [31]. To avoid the problem of low acceptance rates, we resort to using the rejection-free $n$-fold way algorithm as local update [32]. With this computational setup we manage to equilibrate systems with $J_{2}$ as low as $0.23|J_{1}|$ for a range of system sizes which are way out of reach of simple Metropolis simulations. Employing a finite-size scaling (FSS) analysis we determine critical temperatures and exponents, which for all the considered values of $J_{2}$ are compatible with the Onsager exponents of the regular two-dimensional Ising model.
@@
The rest of the paper is organized as follows. The model, simulation details and measured observables are summarized in Sec. II. In Sec. III we introduce the PA algorithm, as well as the adaptations proposed here, and we provide the used simulation parameters. Our results are presented in Sec. IV, and a conclusion is provided in Sec. V.
@@ II.1 Model
where $\sigma_{i}\in\{-1,1\}$ are the spin variables. The first sum is over the nearest-neighbor pairs $\langle ij\rangle$ and the second sum relates to next-nearest neighbors $[ik]$. We choose $J_{1}=1>0$ and $k_{B}=1$ to fix units and consider various $0>J_{2}\in(-1/4,0]$.
Figure 1: Ising model on a honeycomb lattice. Black circles represent spin sites, solid (resp. dashed) lines nearest-neighbor (resp. next-nearest-neighbor) interactions.
The ferromagnetic nn and antiferromagnetic nnn interactions compete with each other, and hence induce frustration, i.e., not all interactions can be satisfied simultaneously. For $J_{2}>-1/4$ the ground state consists of all spins aligned with energy per spin $e_{\mathrm{FM}}=-\frac{3}{2}(1+2J_{2})$. For stronger antiferromagnetic nnn interactions $J_{2}<-1/4$ the model exhibits a peculiar striped ground state with $e_{\mathrm{S}}=-\frac{1}{2}(1-2J_{2})$. This state satisfies two out of three nn and four out of six nnn interactions [18]. In such states certain lines of spins can be flipped at no energetic cost, giving rise to a largely degenerate ground state [19, 28]. Specifically, these states can be mapped to one-dimensional random walks and their degeneracy on a system of edge length $L$ is at least $2^{L}$ [33].
@@ II.2 Observables
To study the critical behavior of the system, we consider various thermodynamic observables. Specifically, from the purely energetic quantities we measure the energy per spin, $e=\langle\mathcal{H}\rangle/N$, and the specific heat, $C_{V}=\left(\langle\mathcal{H}^{2}\rangle-\langle\mathcal{H}\rangle^{2}\right)/NT^{2}$. Here, $T$ denotes the temperature and $N=2L^{2}$ is the number of spins. The purely magnetic quantities considered are the magnetization per spin, $m=\langle|M|\rangle/N=\langle\left|\sum_{l=1}^{N}\sigma_{l}\right|\rangle/N$, and the magnetic susceptibility. Here, we use both the (connected) magnetic susceptibility $\chi=(\langle M^{2}\rangle-\langle|M|\rangle^{2})/NT$, and the high-temperature (disconnected) magnetic susceptibility $\chi^{\prime}=\langle M^{2}\rangle/NT$. Additionally, we consider the mixed quantities,
$\mathrm{d}\hskip 0.50003pt|m|\equiv\frac{\partial}{\partial\beta}\langle|m|\rangle=\langle|m|\rangle\langle\mathcal{H}\rangle-\langle|m|\mathcal{H}\rangle,$ (2)
and
$\mathrm{d}\hskip-0.50003pt\ln|m|\equiv\frac{\partial}{\partial\beta}\ln\langle|m|\rangle=\langle\mathcal{H}\rangle-\frac{\langle|m|\mathcal{H}\rangle}{\langle|m|\rangle},$ (3)
@@
For these quantities we expect the following finite-size scaling (FSS) behavior [34]:
@@
$\mathrm{d}\hskip 0.50003pt|m|_{\max}$ $\propto L^{(1-\beta)/\nu},$ (4e)
$m(\beta_{\max}^{(\mathrm{d}\hskip-0.40974pt\ln|m|)})$ $\propto L^{-\beta/\nu},$ (4f)
$m(\beta_{c})$ $\propto L^{-\beta/\nu},$ (4g)
@@
for different observables $\mathcal{O}$. Note that the correction term with the confluent correction exponent $\omega$ is subleading for any $\omega>1/\nu$. Hence, since $\nu$ is expected to be close to one, and $\omega$ to be larger than one, we drop this term. Subsequently, we obtain the critical exponents $\beta$, $\gamma$, and $\nu$. Equation (4a) represents a logarithmic divergence of the specific heat in the critical regime. Note that this corresponds to the case of the 2D Ising universality class, where $\alpha=0$.
@@ III.1 Population annealing
Population annealing (PA) is an algorithmic framework in which a population of replicas is sequentially cooled while using independent Markov chain Monte Carlo (MCMC) moves at each temperature followed by a population control step [29, 36, 31, 37]. The algorithm can be summarized as follows:
@@
    Initialize a population of $R_{0}=R$ replicas at the initial inverse temperature $\beta_{0}=0$. In some cases a non-zero inverse temperature has to be chosen [38]. Set iteration counter $i\leftarrow 0$.
@@
We choose a population size of R=20  [^0] R=20\,000 throughout. The resampling step 22(b) is realized by the nearest-integer resampling method given by
@@ III.2 Family quantities, correlation, and population annealing performance
Throughout the anneal, the PA performance is monitored by tracking the families of replicas, and related quantities [30]. At $\beta_{i}$, a family refers to a group of replicas that descend from the same replica at $\beta_{0}$. As replicas may be culled during resampling, the number of non-empty families $f$ is a monotonously decreasing function of $\beta$, and the typical size of such a family increases with $\beta$. By construction, replicas from different families are uncorrelated. Thus, the family size indicates the maximal correlation within the population. One way to quantify the size of families is via the replica-averaged family size [30, 40]
@@
where $\mathfrak{N}_{i,k}$ is the number of replicas descending from the initial replica $k$ in the $i$-th annealing step at $\beta=\beta_{i}$. When a PA simulation fails to equilibrate, $\rho_{t}$ typically increases rapidly. Particularly, when $\rho_{t}$ reaches the size of the population, that is when all replicas have the same ancestor and there is only one family left, PA results most likely are no longer reliable.
@@
where the nn blocks are chosen sufficiently large to be effectively uncorrelated, and $\mathcal{O}_{k}^{(n)}$ is the mean of the $k$-th block. A decrease in $R_{\text{eff}}(\mathcal{O})$ may indicate that a PA simulation is falling out of equilibrium. As will be discussed in Sec. III.4, we use $R_{\text{eff}}(E)$ for our adaptive sweep protocol, making sure that this quantity remains almost constant when using this approach.
@@ III.3 Rejection-free PA
Next, we replace the most natural choice of the equilibration routine [step 22(d) in the PA algorithm introduced in Sec. III.1], i.e., the Metropolis algorithm, by the rejection-free $n$-fold way update [32]. For the present model one has $n=(3+1)\times(6+1)\times 2=56$ spin classes (3 nn, 6 nnn, and 2 local spin values). In a first attempt one might be tempted to perform $n_{\text{f}}$ rejection-free updates per site for each replica, implying that each copy ends at a different Metropolis time. As is easily seen, however, this introduces systematic errors particularly at low temperatures  [^1] For a simple way to see this, consider the limit of very small, but non-zero temperatures when effectively only the ground state and the first excited state occur. As there are no available moves with $\Delta E=0$, the rejection-free updates lead to an alternation of ground state and first-excited state configurations, even though in Monte Carlo time the state of lower energy is much more likely.. In contrast, when averaging at a fixed Monte Carlo time (as would be the case when using the Metropolis method) the time step $\Delta t$ is accounted for and the obtained estimates are unbiased. In practice, at first $n_{\text{f}}$ spin flips per site are carried out, such that each replica $k$ arrives at a Metropolis time $t_{k}$. We then determine $\hat{t}=\max_{k}t_{k}$ and run each replica until it reaches $\hat{t}$. This removes the bias, but comes at the cost of reducing the effective level of parallelism as compared to the Metropolis method, as in the worst case all replicas will need to wait for a single replica to reach $\hat{t}$.
@@
After completing this update, observables are measured as ensemble averages as in the original algorithm. At high temperatures this corresponds to of the order of $n_{\text{f}}$ MCS whereas at low temperatures it is equivalent to orders of magnitudes more MCS (see Fig. 6 below).
@@ III.4 Adaptive sweep schedule
Under the assumption that a high degree of decorrelation is linked to good equilibration, we can define a minimal effective population size $R_{\text{eff}}^{\min}$ that the population should have before proceeding to the next temperature. In practice, we repeat the equilibration routine until $R_{\text{eff}}(\mathcal{O})$ is at least $R_{\text{eff}}^{\min}(\beta_{i})=\gamma^{*}R_{i}$ where $\gamma^{*}$ (similar to $\alpha^{*}$) is a constant smaller than one, which we here selected to be $0.9$. If the observable $\mathcal{O}$ was chosen to be the magnetization $M$, then $R_{\text{eff}}(M)$ would (using Metropolis or $n$-fold way updates) almost surely not reach the target effective population size below the critical temperature due to dynamic ergodicity breaking [37]. We hence choose the energy $E$ as the relevant observable $\mathcal{O}$.
Note, that in the above description the number of sweeps necessary is unbounded and thus if the decorrelating effect of the equilibration routine fails to meet the target criterion, the entire simulation will fail to proceed. In a more general setting one should therefore define a cut-off criterion (such as a maximal number of updates or wall-clock time). However, in our case thanks to the combination with rejection-free updates, we did not run into a situation where the adaptive approach spent more time at a temperature than we were willing to allow. Also note that for systems with broken ergodicity this approach is unlikely to be useful, and that for such cases an alternative adaptive sweep protocol using the “restricted autocorrelation time” has been proposed [44].

### v2

# Frustrated Ising model on the honeycomb lattice: Metastability and universality

@@ I Introduction
More recently, the same model on the honeycomb lattice has attracted some attention [18, 19, 20, 21, 22, 23, 24]. Unlike in nonfrustrated spin systems, the lattice geometry plays a vital role for the level of frustration present and therefore affects the macroscopic behavior of the system. This is illustrated by the textbook example of the antiferromagnetic NN Ising model on square and triangular lattices, respectively. The former maps to the ferromagnetic Ising model by a simple transformation, whereas the latter remains disordered at all temperatures and has a nonzero ground-state entropy per spin [25, *Wannier1973Errata]. In fact, precisely this example illustrates the reason why the behavior of the $J_{1}$-$J_{2}$ Ising model is expected to be different on the honeycomb lattice: On the square lattice, the NNN couplings form square lattices, whereas on a honeycomb lattice they form two triangular lattices. Thus, different phenomena may be anticipated for this case, particularly for strong antiferromagnetic NNN interactions.
@@
The character of the ground states in the honeycomb system depends on the relative strength of $J_{1}$ and $J_{2}$ [27]. For $J_{2}>-|J_{1}|/4$, depending on the sign of $J_{1}$, the system either has a ferromagnetically ordered or a Néel-ordered ground state, whereas for $J_{2}<-|J_{1}|/4$ the ground-state manifold is highly degenerate. As is the case on the square lattice, a simple transformation maps negative $J_{1}$ to positive ones (and vice versa). Only quite recently the model was first studied for nonzero temperatures by Bobák et al. [18]. For $J_{2}>-|J_{1}|/4$, using an effective-field theory (EFT) of clusters of different sizes they find evidence of a tricritical point separating a region of second-order transitions for $0.1\lesssim J_{2}/|J_{1}|\leq 0$ from a regime with first-order transitions for $0.25<J_{2}/|J_{1}|\lesssim-0.1$. One of the authors of Ref. [18] subsequently carried out a Monte Carlo (MC) simulation study [20] in which no tricritical point was observed. Instead it showed that the transition remains of second order and within the Ising universality class at least for $J_{2}/|J_{1}|\geq-0.2$. For even lower values of $J_{2}$ simulations showed hysteresis upon cooling and heating owed to a dramatic increase of the autocorrelation times. This increased range of the second-order regime as compared to that of Ref. [18] is supported by a cluster mean-field study [21] which finds only second-order signals for the entire range of $0.25<J_{2}/|J_{1}|\leq 0$. We recently presented some Monte Carlo results of the scaling of the partition function zeros [24], which indicated that the system remains in the Ising universality class at least for $J_{2}/|J_{1}|\geq-0.22$. For $J_{2}/|J_{1}|<-1/4$, on the other hand, it appears unclear whether there is a phase transition at all. Specific-heat peaks are suggestive of some sort of transition [28], and due to the extremely slow relaxation it has been argued that the system shows signs of spin-glass-like ordering [19]. Results of machine learning studies [22, 23] were interpreted as suggestive of two distinct phases seen for small system sizes, hence indicative of the existence of a phase transition.
@@
In the present work, however, we focus on the range in $J_{2}$ with (anti)ferromagnetic order at $T=0$, i.e., $J_{2}/|J_{1}|\in(-1/4,0]$. As one approaches $J_{2}\searrow-0.25|J_{1}|$, standard Monte Carlo techniques such as the Metropolis algorithm are found to be increasingly inefficient due to two effects: First, the introduction of the antiferromagnetic NNN interaction gives rise to local energy minima that simulations tend to get trapped in. Second, as $J_{2}$ is lowered the transition temperature decreases, which translates into low acceptance rates that are further reducing the efficacy of the Metropolis algorithm. For the present study we use the population annealing [29] framework which is well suited for systems with rough free-energy landscapes [30] and which has excellent parallel efficiency [31]. To avoid the problem of low acceptance rates, we resort to using the rejection-free $n$-fold way algorithm as local update [32]. With this computational setup we manage to reliably equilibrate systems with $J_{2}$ as low as $0.23|J_{1}|$ for a range of system sizes which are way out of reach of simple Metropolis simulations. Employing a finite-size scaling (FSS) analysis we determine critical temperatures and exponents, which for all the considered values of $J_{2}$ are compatible with the Onsager exponents of the regular two-dimensional Ising model.
@@
The rest of the paper is organized as follows. The model, simulation details and measured observables are summarized in Sec. II. In Sec. III we introduce the population annealing (PA) algorithm, as well as the adaptations proposed here, and we provide the used simulation parameters. Our results are presented in Sec. IV, and a conclusion is provided in Sec. V.
@@ II.1 Model
where $\sigma_{i}\in\{-1,1\}$ are the spin variables. The first sum is over the nearest-neighbor pairs $\langle ij\rangle$ and the second sum relates to next-nearest neighbors $[ik]$. We choose $J_{1}=1>0$ and $k_{B}=1$ to fix units and consider various negative $J_{2}\in(-1/4,0]$.
Figure 1: Ising model on a honeycomb lattice. Black circles represent spin sites, solid (respectively dashed) lines nearest-neighbor (respectively next-nearest-neighbor) interactions.
The ferromagnetic NN and antiferromagnetic NNN interactions compete with each other, and hence induce frustration, i.e., not all interactions can be satisfied simultaneously. For $J_{2}>-1/4$ the ground state consists of all spins aligned with energy per spin $e_{\mathrm{FM}}=-\frac{3}{2}(1+2J_{2})$. For stronger antiferromagnetic NNN interactions, i.e., for $J_{2}<-1/4$, the model exhibits a peculiar striped ground state with $e_{\mathrm{S}}=-\frac{1}{2}(1-2J_{2})$. This state satisfies two of three NN and four of six NNN interactions [18]. In such states certain lines of spins can be flipped at no energetic cost, giving rise to a largely degenerate ground state [19, 28]. Specifically, these states can be mapped to one-dimensional random walks and their degeneracy on a system of edge length $L$ is at least $2^{L}$ [33].
@@ II.2 Observables
To study the critical behavior of the system, we consider various thermodynamic observables. Specifically, from the purely energetic quantities we measure the energy per spin, $e=\langle\mathcal{H}\rangle/N$, and the specific heat, $C_{V}=\left(\langle\mathcal{H}^{2}\rangle-\langle\mathcal{H}\rangle^{2}\right)/NT^{2}$. Here, $T$ denotes the temperature and $N=2L^{2}$ is the number of spins (see Fig. 13 in Appendix A). The purely magnetic quantities considered are the (absolute) magnetization per spin, $m=\langle|M|\rangle/N$ with $M=\sum_{i=1}^{N}\sigma_{i}$, the fourth-order Binder cumulant $U_{4}=1-\langle M^{4}\rangle/(3\langle M^{2}\rangle^{2})$, and the magnetic susceptibility. Here, we use both the (connected) magnetic susceptibility $\chi=(\langle M^{2}\rangle-\langle|M|\rangle^{2})/NT$, and the high-temperature (disconnected) magnetic susceptibility $\chi^{\prime}=\langle M^{2}\rangle/NT$. Additionally, we consider the mixed quantities,
$\mathrm{d}\hskip 0.50003pt|m|\equiv\frac{1}{N}\frac{\mathrm{d}}{\mathrm{d}\beta}\langle|M|\rangle=\frac{1}{N}\bigl(\langle|M|\rangle\langle\mathcal{H}\rangle-\langle|M|\mathcal{H}\rangle\bigr),$ (2)
$\mathrm{d}\hskip-0.50003pt\ln|m|\equiv\frac{\mathrm{d}}{\mathrm{d}\beta}\ln\langle|M|\rangle=\langle\mathcal{H}\rangle-\frac{\langle|M|\mathcal{H}\rangle}{\langle|M|\rangle},$ (3)
and the derivative of the Binder cumulant
$\frac{\mathrm{d}U_{4}}{\mathrm{d}\beta}=(1-U_{4})\left[\langle\mathcal{H}\rangle-2\frac{\langle M^{2}\mathcal{H}\rangle}{\langle M^{2}\rangle}+\frac{\langle M^{4}\mathcal{H}\rangle}{\langle M^{4}\rangle}\right],$ (4)
@@
For these quantities we expect the following FSS behavior [34]:
@@
$\left[\mathrm{d}U_{4}/\mathrm{d}\beta\right]_{\max}$ $\propto L^{1/\nu},$ (5e)
$\mathrm{d}\hskip 0.50003pt|m|_{\max}$ $\propto L^{(1-\beta)/\nu},$ (5f)
$m(\beta_{\max}^{(\mathrm{d}\hskip-0.40974pt\ln|m|)})$ $\propto L^{-\beta/\nu},$ (5g)
$m(\beta_{c})$ $\propto L^{-\beta/\nu},$ (5h)
@@
for different observables $\mathcal{O}$. Note that the correction term with the confluent correction exponent $\omega$ is subleading for any $\omega>1/\nu$. Hence, since $\nu$ is expected to be close to $1$, and $\omega$ to be larger than $1$, we drop this term. Subsequently, we obtain the critical exponents $\beta$, $\gamma$, and $\nu$. Equation (5a) represents a logarithmic divergence of the specific heat in the critical regime. Note that this corresponds to the case of the two-dimensional Ising universality class, where $\alpha=0$.
@@ III.1 Population annealing
Population annealing is an algorithmic framework in which a population of replicas is sequentially cooled while using independent Markov-chain Monte Carlo (MCMC) moves at each temperature, followed by a population control step [29, 36, 31, 37]. The algorithm can be summarized as follows:
@@
    Initialize a population of $R_{0}=R$ replicas at the initial inverse temperature $\beta_{0}=0$. In some cases a nonzero inverse temperature has to be chosen [38]. Set iteration counter $i\leftarrow 0$.
@@
We choose a population size of R=20  [^0] R=20\,000 throughout. The resampling step 22(b) is realized by the nearest-integer resampling method [30] given by
@@ III.2 Family quantities, correlation, and population annealing performance
Throughout the anneal, the PA performance is monitored by tracking the families of replicas, and related quantities [30]. At $\beta_{i}$, a family refers to a group of replicas that descend from the same replica at $\beta_{0}$. As replicas may be culled during resampling, the number of nonempty families $f$ is a monotonously decreasing function of $\beta$, and the typical size of such a family increases with $\beta$. By construction, replicas from different families are uncorrelated. Thus, the family size indicates the maximal correlation within the population. One way to quantify the size of families is via the replica-averaged family size [30, 40]
@@
where $\mathfrak{N}_{i,k}$ is the number of replicas descending from the initial replica $k$ in the $i$th annealing step at $\beta=\beta_{i}$. When a PA simulation fails to equilibrate, $\rho_{t}$ typically increases rapidly. Particularly, when $\rho_{t}$ reaches the size of the population, that is when all replicas have the same ancestor and there is only one family left, PA results most likely are no longer reliable.
@@
where the $n$ blocks are chosen sufficiently large to be effectively uncorrelated and $\mathcal{O}_{k}^{(n)}$ is the mean of the $k$th block. A decrease in $R_{\text{eff}}(\mathcal{O})$ may indicate that a PA simulation is falling out of equilibrium. As will be discussed in Sec. III.4, we use $R_{\text{eff}}(E)$ for our adaptive sweep protocol, making sure that this quantity remains almost constant when using this approach.
@@ III.3 Rejection-free PA
Next, we replace the most natural choice of the equilibration routine [step 22(d) in the PA framework introduced in Sec. III.1], i.e., the Metropolis algorithm, by the rejection-free $n$-fold way update [32]. For the present model one has $n=(3+1)\times(6+1)\times 2=56$ spin classes (3 NN, 6 NNN, and 2 local spin values). In a first attempt one might be tempted to perform $n_{\text{f}}$ rejection-free updates per site for each replica, implying that each copy ends at a different Metropolis time. As is easily seen, however, this introduces systematic errors particularly at low temperatures  [^1] For a simple way to see this, consider the limit of very small, but nonzero temperatures when effectively only the ground state and the first excited state occur. As there are no available moves with $\Delta E=0$, the rejection-free updates lead to an alternation of ground state and first-excited state configurations, even though in Monte Carlo time the state of lower energy is much more likely.. In contrast, when averaging at a fixed Monte Carlo time (as would be the case when using the Metropolis method) the time step $\Delta t$ is accounted for and the obtained estimates are unbiased. In practice, at first $n_{\text{f}}$ spin flips per site are carried out, such that each replica $k$ arrives at a Metropolis time $t_{k}$. We then determine $\hat{t}=\max_{k}t_{k}$ and run each replica until it reaches $\hat{t}$. This removes the bias, but comes at the cost of reducing the effective level of parallelism as compared to the Metropolis method, as in the worst case all replicas will need to wait for a single replica to reach $\hat{t}$.
@@
After completing this update, observables are estimated as population averages as in the original algorithm. At high temperatures this corresponds to of the order of $n_{\text{f}}$ MCS whereas at low temperatures it is equivalent to orders of magnitudes more MCS (see Fig. 6 in Sec. IV.2 below).
@@ III.4 Adaptive sweep schedule
Under the assumption that a high degree of decorrelation is linked to good equilibration, we can define a minimal effective population size $R_{\text{eff}}^{\min}$ that the population should have before proceeding to the next temperature. In practice, we repeat the equilibration routine until $R_{\text{eff}}(\mathcal{O})$ is at least $R_{\text{eff}}^{\min}(\beta_{i})=\gamma^{*}R_{i}$ where $\gamma^{*}$ (similarly to $\alpha^{*}$) is a constant smaller than $1$, which we here selected to be $0.9$. If the observable $\mathcal{O}$ was chosen to be the magnetization $M$, then $R_{\text{eff}}(M)$ would (using Metropolis or $n$-fold way updates) almost surely not reach the target effective population size below the critical temperature due to dynamic ergodicity breaking [37]. We hence choose the energy $E$ as the relevant observable $\mathcal{O}$.
Note that in the above description the number of sweeps necessary is unbounded and thus if the decorrelating effect of the equilibration routine fails to meet the target criterion, the entire simulation will fail to proceed. In a more general setting one should therefore define a cut-off criterion (such as a maximal number of updates or wall-clock time). However, in our case thanks to the combination with rejection-free updates, we did not run into a situation where the adaptive approach spent more time at a temperature than we were willing to allow. Also note that for systems with broken ergodicity this approach is unlikely to be useful and that for such cases an alternative adaptive sweep protocol using the “restricted autocorrelation time” has been proposed [44].